OpenStudy (anonymous):

lim (sqrt(x+1)-1)/x x->0

OpenStudy (anonymous):

use l hopital

OpenStudy (turingtest):

or if you are not allowed to use l'Hospital's rule, multiply by\[\frac{\sqrt{x+1}+1}{\sqrt{x+1}+1}\]

OpenStudy (anonymous):

what are we multiplying by?

OpenStudy (turingtest):

(sqrt(x+1)+1)/(sqrt(x+1)+1)

OpenStudy (anonymous):

im confused on how you got that